OpenStudy (anonymous):

You invest an initial $1,000 in an account that has an annual interest rate of 2%, compounded quarterly. How much money will you have in the account after 5 years? Round your answer to the nearest whole number.Answer $1,023 $1,094 $1,105 $1,245

OpenStudy (anonymous):

@cherio12

OpenStudy (anonymous):

this is an interest formula one...i dont remember it completely but its along the lines of: \[final money=1000(1+.02)^n\] i'm alittle fuzzy in this area sorry

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

first of all calculate here qurterly rate of interest as 2/4=0.5% (as interest is paid quarterly) and now the no. of periods is 5*4=20 so the amount=1000(1+0.005)^20=1105 (approximately)